# CodeIF-Bench: Evaluating Instruction-Following Capabilities of Large Language Models in Interactive Code Generation

### v3

@@ 6. Threats to Validity
7. Conclusion
This paper presents CodeIF-Bench, the first benchmark for evaluating LLMs’ instruction-following capabilities in interactive code generation. CodeIF-Bench incorporates verifiable instructions aligned with the real-world software development, complemented by corresponding test cases, enabling a comprehensive evaluation of LLMs’ instruction-following performance in multi-turn code generation scenarios. In both Static Conversation and Dynamic Conversation settings, we evaluate the performance of 7 state-of-the-art LLMs and summarize important influencing factors and possible areas for improvement. In our future research, we will expand CodeIF-Bench by incorporating additional programming languages and exploring methods to enhance the interactive code generation capabilities of LLMs.

### v4

@@ 6 Threats to Validity
7 Conclusion
This paper presents CodeIF-Bench, the first benchmark for evaluating LLMs’ instruction-following capabilities in interactive code generation. CodeIF-Bench incorporates verifiable instructions aligned with the real-world software development, complemented by corresponding test cases, enabling a comprehensive evaluation of LLMs’ instruction-following performance in multi-turn code generation scenarios. In both Static Conversation and Dynamic Conversation settings, we evaluate the performance of 6 state-of-the-art LLMs and summarize important influencing factors and possible areas for improvement. In our future research, we will expand CodeIF-Bench by incorporating additional programming languages and exploring methods to enhance the interactive code generation capabilities of LLMs.
Declarations
Funding
This research is supported by the National Natural Science Foundation of China Grants Nos. 62302021 and 62177003.
Author Contributions
Peiding Wang: Data curation, Methodology, Experiments, Writing-original draft; Li Zhang: Methodology, Writing-review & editing; Fang Liu: Data curation, Methodology, Writing-review & editing; Lin Shi: Methodology, Writing-review & editing; Minxiao Li: Data curation, Methodology, Experiments; Bo Shen: Writing-review & editing; An Fu: Writing-review & editing.
Data Availability
The dataset and the source code are available at https://github.com/zhu-zhu-ding/CodeIF-Bench.
Conflict of Interest
The authors have no competing interests to declare that are relevant to the content of this paper.
Ethical Approval
Not applicable to this study.
Informed Consent
Not applicable to this study.